There are 1000 components in a box of which 10 are known to be
defective. two components are selected at random. What is the
probability that
a)both are defective
b)neither is defective
c)just one is defective

Answer by Edwin McCravy(20077)

There are 1000 components in a box of which 10 are known to be
defective. two components are selected at random. What is the
probability that
a)both are defective
Probability that the first one is defective = 10/1000 = 1/100.
Probability that the second one is defective, given that the first
was defective is is 9/999 = 1/111.

P(both are defective) = (10/1000)(9/999) = (1/100)(1/111) = 1/11100

b)neither is defective
There are 990 non-defective ones.

Probability that the first one is not defective = 990/1000 = 99/100.
Probability that the second one is not defective, given that the 
first was not defective is is 989/999.

P(neither is defective) = (99/100)(989/999) = 10879/11100

c)just one is defective
Probability that the first one is defective is 10/1000 or 1/100
Probability that the second one is not defective given that 
the first was defective is 990/999 = 110/111

That's (1/100)(110/111) = 11/1110

Probability that the first one is not defective is 990/1000 or 
99/100
Probability that the second one is defective given that the 
first was not defective is 10/999 

That's also = 11/1110

The probability that one or the other will occur is the sum of 
those two probabilities:

11/1110 + 11/1110 = 11/555

Edwin